OpenStudy (anonymous):

REALLY EASY AND GIVE OUT METALS find angle that is coterminal to 210 and need 2 clockwise

OpenStudy (zpupster):

210 + 360 is positive coterminal =570 210 - 2(360) is th negative coterminal = 210 - 720 = -510